# Are Statistical Methods Obsolete in the Era of Deep Learning?

### v1

@@ 2.3 A note on point estimates
For apples-to-apples comparisons between PINN and MAGI, we will only compare point estimates of the parameters $\bm{\theta}$ and trajectories $\mathbf{X}(t)$, taking the posterior means from MAGI’s Monte Carlo samples. For PINN, following the general practice, we use the neural network’s forward-pass output as its $\mathbf{X}(t)$ estimate, and extract out the corresponding trainable parameters in the neural network architecture when estimating $\bm{\theta}$.
@@ 3 The SEIR model
The SEIR model describes infection progression over time and quantifies disease spread in a population by tracking the flow of individuals through these stages ($S$, $E$, $I$ and $R$):
@@
where $N\equiv S+E+I+R$ is the total population. Three parameters govern the dynamics: the rate of personal contact ($\beta$); the rate of transferring from exposure to infectious ($\sigma$); and the rate of leaving the infectious state ($\gamma$). Despite its relatively simple structure, the SEIR model captures the essential features of many real-world outbreaks and serves as a foundation for more sophisticated or specialized epidemiological models (Li and Muldowney, 1995; Kröger and Schlickeiser, 2020).
@@
One key quantity of interest is the basic reproduction number $R_{0}$, defined as $R_{0}=\beta/\gamma$ (Boonpatcharanon et al., 2022). $R_{0}$ reflects the expected number of secondary infections caused by a single infected individual in a fully susceptible population; it also relates to the (exponential) rate of decrease of the susceptible population: $S(t)=S(0)\exp[-R_{0}(R(t)\,-\,R(0))/N]$. The value of $R_{0}$ serves as a fundamental metric to characterize the transmissibility and provides critical thresholds for public health measures. If $R_{0}>1$, the disease can invade and persist within a population, potentially leading to large-scale outbreaks or epidemics. If $R_{0}<1$ (for example, after intervention), the pathogen’s chain of transmission will be broken, leading to the eventual decline of the disease.
@@ 3.1 Experimental setup
Setting the total population size as $N=100\%$ and using the relationship $S\equiv 100\%-E-I-R$, we can take an equivalent formulation of the SEIR model using three components $E$, $I$, and $R$. Furthermore, for improved numerical stability and because population proportions are nonnegative, for both PINN and MAGI, we work with an equivalent representation of the SEIR system through three log-transformed components, $\log E$, $\log I$, and $\log R$.
@@
The true parameters of the system are set at $\beta=2.0$, $\gamma=0.2$, and $\sigma=0.6$; these values emulate those of highly contagious infectious diseases (such as measles) (Becker and Hasofer, 1998; Paterson et al., 2013; Guerra et al., 2017; Masters et al., 2023). In each simulated dataset, the observation window is from $t=0$ to $t=6$, and a total of 41 equally spaced observations are collected within this time interval. See Figure 1 for an illustration. Observations are simulated with multiplicative log-normal noise at 15%, reflecting realistic measurement variability in epidemiological data. The fully observed case has observations on $E$, $I$ and $R$, whereas the missing-component case only has observations on $I$ and $R$.
In PINN, the hyperparameter $\lambda$ governs the balance between the $L_{2}$ and the physics losses (see Equation (3)). To present the PINN method in the best light, we will show and analyze results for $5$ different values of $\lambda$ at $\lambda=0.1,1,10,100,$ and $1000$.
@@
The PINN configuration and training follow standard recommendations, with van Herten et al. (2022) being one example. More details on PINN configuration are included in the SI Section S1. The MAGI setup, including the default GP kernel and hyperparameter setting, follows the recommended procedure of Yang et al. (2021), which is also detailed in SI Section S2. For both PINN and MAGI, we use a common discretization set $I$ of $161$ evenly-spaced time steps in the observation time period of $t\in[0,6]$ to evaluate the physics-based loss component (see Equation (3)) and $W_{I}$ (see Equation (4), respectively. For forecasting, an additional $160$ discretization time steps are used for the out-of-sample period of $t\in(6,12]$. Python codes and the simulated datasets are deposited in an anonymous GitHub repository [^1] Code for reproducing all results can be found at https://anonymous.4open.science/r/stat-vs-dl-71B2/..
@@ 3.2 Results from the fully observed case
Figure 1: Trajectory reconstruction and prediction by PINN and MAGI for the SEIR model in the fully observed case. The dots show one sample dataset (out of 100). The dashed black lines give the true curves, which are to be identified. Each solid blue curve is the estimate from one dataset. The red dashed vertical line separates the in-sample observation period from the future forecasting period. Top row: MAGI estimates. Lower two rows: PINN estimates, with $\lambda=10$ and $\lambda=1000$, respectively. Each solid blue curve is the estimate from one dataset.
@@
It is evident from Figure 1 that the performance of PINN is highly dependent on its hyperparameter $\lambda$: $\lambda=10$ significantly outperforms $\lambda=1000$. SI Figure S2 further corroborates this observation by examining a broader range of $\lambda$ values. Moreover, even with the best $\lambda=10$ setting, PINN occasionally produces outlier trajectories that do not fit the observed data. PINN’s tendency to produce outliers is also seen in Figure 2, which contains boxplots of in-sample trajectory reconstruction error (Equation (9)) and out-of-sample forecasting error (Equation (10)). Given that PINN is trained on a large number (60K) of epochs, such outliers are likely due to the optimizer becoming trapped in local modes. Despite PINN’s occasional outliers and greater sensitivity to the choice of the hyperparameter $\lambda$, the majority of well-behaved in-sample trajectory reconstruction instances of PINN and MAGI are roughly comparable, as shown in the top row of Figure 2. However, when we move on to forecasting the system at future times, PINN becomes notably worse than MAGI, as shown in the bottom row of Figure 2. One possible explanation is that, without data to guide PINN at future time points, it struggles to optimize the over-parameterized network to satisfy the ODE. On the contrary, the GP used by MAGI is much more lightweight, making it easier to explore the sampling space and converge towards the correct solution.
@@
Figure 2: Boxplots showing the RMSE on the logarithm of the SEIR system components across 100 datasets in the fully observed case. Lower value indicates better performance. The y-axis is displayed in the logarithmic scale for better visualization. Top row: in-sample trajectory reconstruction (Equation (9)); bottom row: future forecasting (Equation (10)). The three columns correspond to three system components: $E$, $I$, and $R$. In each panel, the leftmost boxplot is for MAGI, and the remaining boxplots are for PINN under different hyperparameters $\lambda$; the dashed vertical line separates MAGI and PINN.
@@
The parameter estimation results offer a complementary perspective. Figure 3 shows the boxplots of the absolute estimating errors for each of the parameters, together with those for the three key quantities of interest. Despite less accurate trajectory forecasting as illustrated in Figure 1, PINN is still able to recover key parameters ($\beta$ and $R_{0}$) competitively to MAGI. This is because, first, the in-sample fittings of PINN and MAGI are comparable, which is more important for parameter estimation. Second, it is well understood from the literature on two-stage approaches for inverse problems that trajectory reconstruction does not have to be perfect in order to recover the ODE parameters (Bär et al., 1999; Müller and Timmer, 2004; Rai and Tripathi, 2019). The estimated trajectory can be thought of as the response variable, and estimating the ODE parameter can be thought of as carrying out a regression, which could lead to reasonable parameter estimation even though there is noise in the response variable. However, it is also well known that such correspondence becomes less reliable when there are missing components. Indeed, as we will see, the performance of PINN will deteriorate in the missing component case.
@@
Figure 3: Boxplots showing the absolute errors in parameter estimation for PINN and MAGI across 100 datasets in the fully observed case. Lower value indicates better performance. The y-axis is displayed in the logarithmic scale for better visualization. Top row: The errors for the original parameters $\beta$, $\gamma$, and $\sigma$. Bottom row: The errors for $R_{0}$, peak timing, and peak intensity – our three quantities of interest. In each panel, the leftmost blue boxplot is for MAGI, and the remaining red boxplots are for PINN under different $\lambda$ settings; the dashed vertical line separates MAGI and PINN results.
@@ 3.3 Results from the missing component case
Figure 4: Trajectory reconstruction and prediction by PINN and MAGI for the SEIR model in the missing component case, where the $E$ component is unobserved. The dots show one sample dataset (out of 100). The legend and layout of this figure are identical to Figure 1; see the caption there.
@@
As shown in Figure 4, the MAGI method fits the in-sample $I(t)$ and $R(t)$ accurately while recovering a plausible range of $E(t)$ from the observed data. The GP prior in MAGI provides shrinkage for the $E$ component, effectively reducing the variance in its recovery. In contrast, PINN exhibits even greater sensitivity to the hyperparameter $\lambda$ and produces more noticeable outliers, as shown in Figure 4, and in SI Figure S3 for a wider range of $\lambda$.
Figure 5: Boxplots showing the RMSE on the logarithm of the SEIR system components across 100 datasets in the missing $E$ component case. The legend and layout of this figure are identical to Figure 2; see the caption there.
@@
Figure 6: Boxplots showing the absolute errors in parameter estimation for PINN and MAGI across 100 datasets in the missing $E$ component case. The legend and layout of this figure are identical to Figure 3; see the caption there.
@@
MAGI’s performance results are not without limitations. As a Bayesian approach, MAGI is subject to prior distribution-induced bias. As shown in Figure 4, MAGI’s inferred trajectories of the missing component $E$ tend to be biased higher than the ground truth, while the future predictions for $I$ and $R$ are slightly lower. This bias could come from the GP prior, which favors smooth curves around average in-sample values. Interestingly, a similar bias is also observed in the PINN results. Additionally, PINN’s occasional outlier predictions resemble a flat line. This phenomenon could be attributable to the NN shrinking toward its initial weights during training, with random initializations prone to outputting flat lines due to variance decay across network layers (Glorot and Bengio, 2010).
@@ 4 The Lorenz model
In this section, we compare PINN and MAGI on the Lorenz system, one of the most well-known chaotic dynamical systems (Lorenz, 1963; Hirsch et al., 2012; Sparrow, 2012). Intuitively, a chaotic system is one whose trajectories are extremely sensitive to perturbations in parameters and initial conditions (Strogatz, 2018). The Lorenz system has three components $(X,Y,Z)$ governed by the following ODEs, parameterized by $\beta$, $\rho$, and $\sigma$:
@@
The Lorenz system is a good test case due to its chaotic potential, numerical pathologies, and unpredictable behavior. For example, while some Lorenz trajectories will converge to a stationary point (i.e., stable behavior), changing the parameters slightly will yield trajectories that oscillate in the state space in a characteristic butterfly-shaped path for perpetuity without convergence (Strogatz, 2018; Sparrow, 2012).
We choose a “Chaotic (Butterfly)” setting to compare MAGI and PINN. Figure 7 depicts the classic chaotic butterfly pattern of the Lorenz system, which, in this case, oscillates in perpetuity. This regime is mathematically chaotic (Hirsch et al., 2012), with parameters $\bm{\theta}=(\beta,\rho,\sigma)=(\frac{8}{3},28,10)$ and initial conditions $(X(0),Y(0),Z(0))=(5,5,5)$. We generate 100 independent datasets, and then apply PINN and MAGI to each dataset. We consider parameter estimation, in-sample trajectory reconstruction and the prediction of future trajectories in the comparison; the latter two are measured by RMSE and $\text{RMSE}_{\text{pred}}$ in Equations (9) and (10) respectively. Additionally, we also analyzed one stable regime of the Lorenz system and report the results in SI Section S5.1.

### v2

@@ 2.4 A note on point estimates and uncertainty quantification
for fair comparisons between PINN and MAGI, only compare point estimates of the parameters $\bm{\theta}$ and trajectories $\mathbf{X}(t)$, taking the posterior means from MAGI’s Monte Carlo samples. For PINN, following the general practice, we use the neural network’s forward-pass output as its $\mathbf{X}(t)$ estimate, and extract out the corresponding trainable parameters in the neural network architecture when estimating $\bm{\theta}$.
@@ 3 The SEIR model
The SEIR model describes infection progression over time and quantifies disease spread in a population by tracking the flow of individuals through these stages ($S$, $E$, II and $R$):
@@
where $N\equiv S+E+I+R$ is the total population. Three parameters govern the dynamics: the rate of personal contact ($\beta$); the rate of transferring from exposure to infectious ($\sigma$); and the rate of leaving the infectious state ($\gamma$). Despite its relatively simple structure, the SEIR model captures the essential features of many real-world outbreaks and serves as a foundation for more sophisticated or specialized epidemiological models (Li and Muldowney 1995, Kröger and Schlickeiser 2020).
@@
One key quantity of interest is the basic reproduction number $R_{0}$, defined as $R_{0}=\beta/\gamma$ (Boonpatcharanon et al. 2022). $R_{0}$ reflects the expected number of secondary infections caused by a single infected individual in a fully susceptible population; it also relates to the (exponential) rate of decrease of the susceptible population: $S(t)=S(0)\exp[-R_{0}(R(t)\,-\,R(0))/N]$. The value of $R_{0}$ serves as a fundamental metric to characterize the transmissibility and provides critical thresholds for public health measures. If $R_{0}>1$, the disease can invade and persist within a population, potentially leading to large-scale outbreaks or epidemics. If $R_{0}<1$, the pathogen’s chain of transmission will be broken, leading to the eventual decline of the disease.
@@ 3.1 Experimental setup
Setting the total population size as $N=100\%$ and using the relationship $S\equiv 100\%-E-I-R$, we can take an equivalent formulation of the SEIR model using three components $E$, II, and $R$. Furthermore, for improved numerical stability and because population proportions are nonnegative, for both PINN and MAGI, we work with an equivalent representation of the SEIR system through three log-transformed components, $\log E$, $\log I$, and $\log R$.
@@
The true parameters of the system are set at $\beta=2.0$, $\gamma=0.2$, and $\sigma=0.6$; these values emulate those of highly contagious infectious diseases (such as measles) (Becker and Hasofer 1998, Paterson et al. 2013, Guerra et al. 2017, Masters et al. 2023). In each simulated dataset, the observation window is from $t=0$ to $t=6$, and a total of 41 equally spaced observations are collected within this time interval. See Figure 1 for an illustration. Observations are simulated with multiplicative log-normal noise at 15%, reflecting realistic measurement variability in epidemiological data. The fully observed case has observations on $E$, II and $R$, whereas the missing-component case only has observations on $I$ and $R$.
@@
For both PINN and MAGI, we use a common discretization set of $161$ evenly-spaced time steps in the observation time period of $t\in[0,6]$ to evaluate the physics-based loss component (see Equation (3)) and $W_{I}$ (see Equation (4)), respectively. For forecasting, an additional $160$ discretization time steps are used for the out-of-sample period of $t\in(6,12]$. Python codes and the simulated datasets are deposited in an anonymous GitHub repository [^1] Code for reproducing all results can be found at https://anonymous.4open.science/r/tas-revisions-7650/..
3.1.1 PINN implementation details
In PINN, the hyperparameter $\lambda$ governs the balance between the $L_{2}$ and the physics losses (see Equation (3)). Rather than relying on a single “out-of-the-box” PINN setting, we use established recommendations as baselines and systematically sweep the main tunable choices to assess transferability and report best-case performance. We will show and analyze results for $5$ different values of $\lambda$ at $\lambda=0.1,1,10,100,$ and $1000$. We also consider two PINN implementations: one based on a practitioner codebase in the PINN literature (van Herten et al. 2022), and one using the widely-adopted DeepXDE package (Lu et al. 2021). Within the DeepXDE framework, we adopt the recommended architecture following the DeepXDE documentations (Lu et al. 2021) and examples (Wang et al. 2022). In addition to the $5$ values of $\lambda$, we vary (i) $2$ network layer-size configurations ($3$ hidden layers of $40$ or $512$ units); (ii) $2$ learning rate scheduler choices (constant or exponential decay); and (iii) whether to fine-tune with second-order L-BFGS after first-order Adam training (Rathore et al. 2024). Additional PINN details are in SI Section S1. For each of the resulting $5\times 2\times 2\times 2=40$ configurations, we run $5$ seeds and select the best-performing configuration per $\lambda$ value based on inverse problem accuracy for in-sample experiments, or forecast accuracy for forecasting problems. This post-hoc selection gives PINN an explicit advantage: if MAGI still outperforms the best-case PINN, the conclusion is only strengthened, since it would outperform any other configuration as well. We report these best-case PINN results in the main text and defer comprehensive comparisons across other configurations to SI.
3.1.2 MAGI implementation details
The MAGI setup follows the recommended procedure of Yang et al. (2021) and is also detailed in SI Section S2. Unlike PINN, MAGI does not require sweeping over hyperparameter configurations: the Gaussian process hyperparameters are estimated in a statistically principled manner by maximizing the marginal likelihood of the GP smoothing of the observed data. This automatic procedure is used in all experiments except the SEIR missing-component case, where the entirely unobserved $E$ component lacks data for marginal likelihood estimation; in that setting, the Matérn kernel hyperparameters $(\phi_{1},\phi_{2})$ for $E$ are manually specified, guided by the interpolation and optimization routine described in SI Section S2. We note that across our PINN configuration grid, the missing-component setting is similarly challenging: neither method produces reliable results under purely automatic tuning.
For the fully observed setting, we perform (a) $1,000$ burn-in and $1,000$ sampling steps for in-sample inference, followed by (b) $1,000$ burn-in and $1,000$ sampling steps for forecasting. In the missing component setting, we perform (a) $10,000$ burn-in and $5,000$ sampling steps for in-sample inference, followed by (b) $5,000$ burn-in and $100,000$ sampling steps for forecasting. These additional steps encourage more thorough exploration of the less-identifiable parameter space.
For numerical stability, we add a small diagonal nugget (jitter) term $10^{-6}$ to GP covariance matrices to ensure numerical stability and invertibility in Cholesky-based computations, as is standard in GP toolkits (e.g., GPyTorch). This nugget is fixed for numerical stability and is not treated as a tunable hyperparameter.
@@ 3.2 Results from the fully observed case
Figure 1: Trajectory reconstruction and prediction by best-case PINN and MAGI for the SEIR model in the fully observed case. The dots show one sample dataset (out of 100). The dashed black lines give the true curves, which are to be identified. Each solid blue curve is the estimate from one dataset. The red dashed vertical line separates the in-sample observation period from the future forecasting period. Top row: MAGI estimates. Lower two rows: best-case PINN estimates, with $\lambda=10$ and $\lambda=1000$, respectively. Each solid blue curve is the estimate from one dataset. $S(t)$ is not plotted because it is completely determined once $E(t),I(t)$ and $R(t)$ are known: $S\equiv 100\%-E-I-R$.
@@
It is evident from Figure 1 that the performance of PINN, even under its best-case configuration, is highly sensitive to the choice of the hyperparameter $\lambda$: in particular, $\lambda=10$ substantially outperforms $\lambda=1000$. Section S4 in the Supplementary Information further corroborates this observation by examining a broader range of $\lambda$ values as well as other PINN configurations. In the remainder of the paper, we focus on the best-case PINN and, whenever it is clear from context, refer to it simply as PINN. Upon closer examination, even with the best $\lambda=10$ setting, PINN occasionally produces outlier trajectories that do not fit the observed data. PINN’s tendency to produce outliers is also seen in Figure 2, which contains boxplots of in-sample trajectory reconstruction error (Equation (9)) and out-of-sample forecasting error (Equation (10)). Given that PINN is trained on a large number (60K or 300K) of epochs, such outliers are likely due to the optimizer becoming trapped in local modes. Despite PINN’s occasional outliers and greater sensitivity to the choice of the hyperparameter $\lambda$, the majority of well-behaved in-sample trajectory reconstruction instances of PINN and MAGI are roughly comparable, as shown in the top row of Figure 2. However, when we move on to forecasting the system at future times, PINN becomes notably worse than MAGI, as shown in the bottom row of Figure 2. One possible explanation is that, without data to guide PINN at future time points, it struggles to optimize the over-parameterized network to satisfy the ODE. On the contrary, the GP used by MAGI is much more lightweight, making it easier to explore the sampling space and converge towards the correct solution.
@@
Figure 2: Boxplots showing the RMSE on the logarithm of the SEIR system components across 100 datasets in the fully observed case. Lower value indicates better performance. The y-axis is displayed in the logarithmic scale for better visualization. Top row: in-sample trajectory reconstruction (Equation (9)); bottom row: future forecasting (Equation (10)). The three columns correspond to three system components: $E$, II, and $R$. In each panel, the leftmost boxplot is for MAGI, and the remaining boxplots are for best-case PINN under different hyperparameters $\lambda$; the dashed vertical line separates MAGI and best-case PINN. Comparison with other PINN configurations can be found in SI Figure S3.
@@
The parameter estimation results offer a complementary perspective. Figure 3 shows the boxplots of the absolute estimating errors for each of the parameters, together with those for the three key quantities of interest. Despite less accurate trajectory forecasting as illustrated in Figure 1, PINN is still able to recover key parameters ($\beta$ and $R_{0}$) competitively to MAGI. This is because, first, the in-sample fittings of PINN and MAGI are comparable, which is more important for parameter estimation. Second, it is well understood from the literature on two-stage approaches for inverse problems that trajectory reconstruction does not have to be perfect in order to recover the ODE parameters (Bär et al. 1999, Müller and Timmer 2004, Rai and Tripathi 2019). The estimated trajectory can be thought of as the response variable, and estimating the ODE parameter can be thought of as carrying out a regression, which could lead to reasonable parameter estimation even though there is noise in the response variable. However, it is also well known that such correspondence becomes less reliable when there are missing components. Indeed, as we will see, the performance of PINN will deteriorate in the missing component case.
@@
Figure 3: Boxplots showing the absolute errors in parameter estimation for best-case PINN and MAGI across 100 datasets in the fully observed case. Lower value indicates better performance. The y-axis is displayed in the logarithmic scale for better visualization. Top row: The errors for the original parameters $\beta$, $\gamma$, and $\sigma$. Bottom row: The errors for $R_{0}$, peak timing, and peak intensity – our three quantities of interest. In each panel, the leftmost blue boxplot is for MAGI, and the remaining red boxplots are for best-case PINN under different $\lambda$ settings; the dashed vertical line separates MAGI and PINN results. Comparison with other PINN configurations can be found in SI Figure S4.
@@ 3.3 Results from the missing component case
Figure 4: Trajectory reconstruction and prediction by best-case PINN and MAGI for the SEIR model in the missing component case, where the $E$ component is unobserved. The dots show one sample dataset (out of 100). The legend and layout of this figure are identical to Figure 1; see the caption there.
@@
As shown in Figure 4, the MAGI method fits the in-sample $I(t)$ and $R(t)$ accurately while recovering a plausible range of $E(t)$ from the observed data. The GP prior in MAGI provides stability for the $E$ component, effectively reducing the variance in its recovery. In contrast, PINN exhibits even greater sensitivity to the hyperparameter $\lambda$ and produces more noticeable outliers, as shown in Figure 4, and in SI Figure S5 for a wider range of $\lambda$.
Figure 5: Boxplots showing the RMSE on the logarithm of the SEIR system components across 100 datasets in the missing $E$ component case. The legend and layout of this figure are identical to Figure 2; see the caption there. Only best-case PINN configurations are reported here. Results for other PINN configurations are in SI Figure S6.
@@
Figure 6: Boxplots showing the absolute errors in parameter estimation for PINN and MAGI across 100 datasets in the missing $E$ component case. The legend and layout of this figure are identical to Figure 3; see the caption there. Only best-case PINN configurations are reported here. Results for other inferior PINN configurations are in SI Figure S7.
@@
MAGI’s performance results are not without limitations. As a Bayesian approach, MAGI is subject to prior-distribution-induced bias. As shown in Figure 4, MAGI’s inferred trajectories of the missing component $E$ tend to be higher than the ground truth, while the future predictions for $I$ and $R$ are slightly lower. This could come from the GP prior, which favors smooth curves around average in-sample values. Interestingly, a similar bias is also observed in the PINN results. Additionally, PINN’s occasional outlier predictions resemble a flat line. This phenomenon could be attributable to the NN shrinking toward its initial weights during training, with random initializations prone to outputting flat lines due to variance decay across network layers (Glorot and Bengio 2010).
@@ 4 The Lorenz model
In this section, we compare PINN and MAGI on the Lorenz system, one of the most well-known chaotic dynamical systems (Lorenz 1963, Hirsch et al. 2012, Sparrow 2012). Intuitively, a chaotic system is one whose trajectories are extremely sensitive to perturbations in parameters and initial conditions (Strogatz 2018). The Lorenz system has three components $(X,Y,Z)$ governed by the following ODEs, parameterized by $\beta$, $\rho$, and $\sigma$:
@@
The Lorenz system is a good test case due to its chaotic potential, numerical pathologies, and unpredictable behavior. For example, while some Lorenz trajectories will converge to a stationary point (i.e., stable behavior), changing the parameters slightly will yield trajectories that oscillate in the state space in a characteristic butterfly-shaped path for perpetuity without convergence (Strogatz 2018, Sparrow 2012).
We choose a “Chaotic (Butterfly)” setting to compare MAGI and PINN. Figure 7 depicts the classic chaotic butterfly pattern of the Lorenz system, which, in this case, oscillates in perpetuity. This regime is mathematically chaotic (Hirsch et al. 2012), with parameters $\bm{\theta}=(\beta,\rho,\sigma)=(\frac{8}{3},28,10)$ and initial conditions $(X(0),Y(0),Z(0))=(5,5,5)$. We generate 100 independent datasets, and then apply PINN and MAGI to each dataset. We consider parameter estimation, in-sample trajectory reconstruction and the prediction of future trajectories in the comparison; the latter two are measured by RMSE and $\text{RMSE}_{\text{pred}}$ in Equations (9) and (10) respectively. Additionally, we also analyzed one stable regime of the Lorenz system and report the results in SI Section S6.